OpenStudy (ooxbrittanyxoo):

help will medal Arrange the equations in the correct sequence to rewrite the formula for displacement. d=1/2(v0+v)t. to find v0. In the formula, d is displacement, v0 is the initial velocity, v is the final velocity, and t is the time.

OpenStudy (tkhunny):

Have you considered multiplying by 2 to get you on your way?

OpenStudy (ooxbrittanyxoo):

TILES A) \[2d \div t=v0+v\] B) \[v0=2d \div t+v\] C) \[2d=(v0+v)t\] D) \[2t \div d=v0+v\] E) \[v0=2d \div t-v\] SEQUENCE [box1] v [box2] v [box3]

OpenStudy (ooxbrittanyxoo):

so would C) be the first one?

OpenStudy (ooxbrittanyxoo):

would it be C,A,E?

OpenStudy (tkhunny):

What did you get after multiplying by 2 and dividing by t?